OpenStudy (anonymous):

Hi I need some help with this problem. The problem statement, all variables and given/known data it reads this way: The manufacturer of a digital ammeter quotes its uncertainty as ± 1.5% ± 2 digits. a) Determine the uncertainty in a constant reading of 2.64A b) The meter is used to measure the current from a d.c power supply. The current is found to fluctuate randomly between 1.58A and 2.04A. Determine the most likely value of the current, with its uncertainty. ..,Part (a) is ok,, my answer is right (answrs on back of the book)...but part (b) was just too hard..answr given: 2.01±0.09A

OpenStudy (anonymous):

a) find 1.5% of 2.64 and give in 2 significant figures - (2.64/100) x 1.5 = 0.0396 so uncertainty is 2.64 ±0.039. b) ±, thinking about that I think, maximum + on lowest ampere, and maximum minus on highest ampere, so 1.58/100 * 1.5 = 0.0237 & 2.04/100 * 1.5 = 0.0306. so 1.58A ==> 1.603A & 2.04A ===> 2.0094A difference is 2.0094A - 1.603A = 0.4062A SO: 1.58A + 0.4026A = 1.9826A and 2.04A - 0.4026A = 1.6374A median of these two, i.e the central point, would be the mean, so (1.9826 + 1.6374)/2 = 1.81A and the uncertainty would be ± 0.09A so 1.81 ± 0.09A, is what I got. But I'm bad with errors, thats what I THINK. If anyone could blow us out the water that would be great.